# Individualised Treatment Effects Estimation with Composite Treatments and Composite Outcomes

### v1

@@ I Introduction
Causal inference helps to answer fundamental “what if” questions regarding the effect of a set of variables on a set of outcome variables of interest. For instance, one might be interested in estimating the risk of cancer for an individual if they were to quit smoking. Such questions are prevalent in science, including healthcare, economics, education, social science, marketing, and computer science. While randomised controlled trials remain the gold standard for causal effect estimation, they can be time-consuming, expensive, ethically challenging, and more importantly not suitable for individual-level causal effect estimation due to their strict inclusion criteria and small and less representative samples. By contrast, non-experimental approaches based on observational datasets (e.g., electronic health records) address many of these limitations and can complement randomised controlled trials. Recently, there has been notable progress in causal machine learning [2, 8], a discipline that integrates rigorous causal inference methodologies with state-of-the-art machine learning techniques, particularly for the analysis of observational data.
Existing research in individualised treatment effects (ITE), also referred to as conditional average treatment effects estimation or heterogeneous treatment effects estimation in the machine learning literature [39], is mostly focused on simple settings considering single treatments, such as binary treatments, and single outcomes [13, 6, 20, 10, 30]. However, the real world is complex and might involve composite treatments, i.e., multiple simultaneous treatments (often referred to as polypharmacy in clinical literature [17]), as well as composite outcomes, where we aim to optimise multiple simultaneous outcomes of interest. Thus, existing methods are not always suitable for practical adoption in these multi-faceted scenarios.
@@
Motivating example: Suppose a patient is admitted to a hospital for heart surgery. Approximately 20–50% of such patients are at risk of atrial fibrillation within the first seven days after surgery [15], which is associated with an increased risk of heart failure, dementia, and stroke [31]. During the surgery, clinicians can choose from more than ten interventions (e.g., beta-blockers, statins, and others) to manage multiple outcomes, including atrial fibrillation, mortality, and readmission to the intensive care unit, as depicted in Fig. 1. Evaluating the combined effects of these interventions on the outcomes of interest is a complex open problem that urgently requires solutions with potentially significant impact on healthcare.
@@ II-A Individualised Treatment Effects Estimation
After the pioneering work of [23], a wide variety of machine learning-based ITE learners have been proposed [2]. These methods can be broadly categorised as: (i) representation learning or deep learning-based learners, e.g., [13, 39, 18, 19, 6, 8], (ii) tree-based learners [1], and (iii) meta-learners, which are model-agnostic and can be subdivided into direct learners (S-Learner and T-Learner) [26], and indirect learners (RA-Learner, DR-Learner, and X-Learner) [24, 12, 26]. For a comprehensive review, please refer to [11]. While these frameworks have demonstrated success, they predominantly address settings with a single treatment and a single outcome.
The study of composite treatment effects has been constrained by data scarcity arising from the exponential number of treatment combinations (for $K$ binary treatments, $2^{K}$ combinations) and the possibility that many combinations have insufficient or no data at all. In this context, [45] proposed an S-Learner-based approach, called deconfounder (DEC), that accommodates composite (binary and real) treatments for a single outcome. They leverage proxy variables and latent representations to capture interactions among multiple treatments and mitigate unmeasured confounding. Similarly, [50] developed the variational sample re-weighting (VSR) approach—an extension of S-Learner—to handle high-dimensional binary treatments (referred to as bundle treatments), using learned latent representations to de-correlate treatments from confounders. To address data scarcity, [34] proposed a data-augmentation strategy for composite treatment and single outcome settings, whereby $K$ ITE learners generate balanced datasets for each treatment before training a standard supervised deep learning model to predict potential outcomes. However, this method requires to train multiple models and is also limited to binary composite treatments.
@@ III Background
Notations: Let $\mathbf{x}_{i}=(x_{i1},\dots,x_{ip})^{\top}$ denote the $p$-dimensional vector of pre-treatment features for patient $i$ ($i=1,\dots,N$). Let $\mathbf{t}_{i}=(t_{i1},\dots,t_{iK})^{\top}$ be the $K$-dimensional treatment vector for patient $i$, where $t_{ik}$ represents the $k$-th component of the potentially composite treatment. Let $\mathbf{y}_{i}=(y_{i1},\dots,y_{iM})^{\top}$ be the $M$-dimensional outcome vector for patient $i$, where $y_{im}$ is the $m$-th outcome. $y_{im}(\mathbf{t}_{i})$ denotes the potential outcome for patient $i$ on outcome $m$ if they were to receive treatment vector $\mathbf{t}$. The observed outcome is $y_{im}=y_{im}(\mathbf{t}_{i})$. The observational dataset is $\mathcal{D}=\{(\mathbf{x}_{i},\mathbf{t}_{i},\mathbf{y}_{i})\}_{i=1}^{N}$.
@@ IV Methods
Motivation: Complex real-world scenarios often entail multiple concurrent treatments (polypharmacy) and multiple outcomes of interest. Existing research on ITE estimation, however, typically focuses on single treatment and single outcome problems. In the composite treatment and composite outcome setting, the exponential number of possible treatment combinations (e.g., $2^{K}$ for $K$ binary treatments) leads to data scarcity, where certain treatment combinations may have limited or no samples. Such data sparsity poses a significant challenge to reliably estimating ITE.
@@
Figure 2: Architecture of H-Learner. It comprises two networks: (1) a hypernetwork conditioned on treatments and outcomes, and (2) a target network (ITE learner). The hypernetwork generates the target network’s weights for each treatment-outcome combination, enabling dynamic information sharing.
@@
Figure 4: Comparative study of ITE learners for binary composite treatments and composite outcomes (shaded region represents one standard error over 10 repetitions). H-Learner does not need any assumptions about the interaction among the treatments and can handle composite outcomes but SCP needs to know the causal structure among treatments and is limited to single outcomes.
Neural Architecture and Working: Figure 2 illustrates the key components of H-Learner, which include an embedding layer, a hypernetwork, and a target network (the ITE learner). The embedding layer learns representations for each of the unique combinations of treatments and outcomes and has been implemented using a single linear layer. The embedding layer maps given causes and an effect to a meaningful representation which is equivalent to learning embeddings for words in natural language processing [44].
@@
Training proceeds by feeding the treatments and outcomes into the embedding layer, whose outputs are passed to the hypernetwork. The hypernetwork produces weights for the target network, which processes the observed covariates to predict potential outcomes. We compute the prediction loss on factual observations (the data points for which the treatments and outcomes are known) and backpropagate the error. Since only the embedding layer and the hypernetwork parameters are trainable, the backpropagated gradients update those components to learn better embeddings for treatments and outcomes and improve the hypernetwork’s mapping from embeddings to target-network weights. Each training epoch thus leads to dynamic information sharing across the various treatments and outcomes, benefiting low-data configurations through mechanisms similar to transfer learning and multitasking [29, 7].
@@ V-B Results
In this subsection, we compare H-Learner under two scenarios: (1) arbitrary composite treatments and (2) binary composite treatments.
@@ VI Conclusion
Statements and Declarations
Competing Interests
The authors declare that they have no competing interests.
Code Availability
The finalised code, after acceptance, will be available at https://github.com/jmdvinodjmd/HLearner.

### v2

@@ Abstract
Clinical Relevance—This paper develops a novel methodology, the H-Learner, to estimate individualised treatment effects when clinicians need to consider multiple interventions simultaneously to impact several patient outcomes. By accurately estimating these complex, individualised effects from observational data, this method has the potential to provide clinicians with more precise insights for tailoring treatment strategies, ultimately leading to improved patient care and outcomes in complex clinical scenarios where single interventions and outcomes are insufficient to capture the full picture.
@@ I INTRODUCTION
Causal inference [^1] This work is accepted in the 47th Annual International Conference of the IEEE Engineering in Medicine and Biology Society, to be held in Bella Center, Copenhagen, Denmark, July 14-17, 2025. helps to answer fundamental “what if” questions regarding the effect of a set of variables on a set of outcome variables of interest. For instance, one might be interested in estimating the risk of cancer for an individual if they were to quit smoking. Such questions are prevalent in science, including healthcare, economics, education, social science, marketing, and computer science. While randomised controlled trials remain the gold standard for causal effect estimation, they can be time-consuming, expensive, ethically challenging, and more importantly not suitable for individual-level causal effect estimation due to their strict inclusion criteria and small and less representative samples. By contrast, non-experimental approaches based on observational datasets (e.g., electronic health records) address many of these limitations and can complement randomised controlled trials. Recently, there has been notable progress in causal machine learning [2, 7], a discipline that integrates rigorous causal inference methodologies with state-of-the-art machine learning techniques, particularly for the analysis of observational data.
Existing research in individualised treatment effects (ITE), also referred to as conditional average treatment effects estimation or heterogeneous treatment effects estimation in the machine learning literature [38], is mostly focused on simple settings considering single treatments, such as binary treatments, and single outcomes [12, 6, 19, 9, 29]. However, the real world is complex and might involve composite treatments, i.e., multiple simultaneous treatments (often referred to as polypharmacy in clinical literature in the context of medications [16], or multimodal treatment more broadly), as well as composite outcomes, where we aim to optimise multiple simultaneous outcomes of interest. Thus, existing methods are not always suitable for practical adoption in these multi-faceted scenarios.
@@
Motivating example: Suppose a patient is admitted to a hospital for heart surgery. Approximately 20–50% of such patients are at risk of atrial fibrillation within the first seven days after surgery [14], which is associated with an increased risk of heart failure, dementia, and stroke [30]. During and after the surgery, clinicians can choose from multiple interventions (e.g., beta-blockers, statins, DC cardioversion etc) to manage multiple outcomes, including atrial fibrillation, mortality, and readmission to the intensive care unit, as depicted in Fig. 1. Evaluating the combined effects of these interventions on the outcomes of interest is a complex open problem that urgently requires solutions with potentially significant impact on healthcare.
@@ II-A Individualised Treatment Effects Estimation
After the pioneering work of Johansson et al. [22], a wide variety of machine learning-based ITE learners have been proposed [2]. These methods can be broadly categorised as: (i) representation learning or deep learning-based learners, e.g., [12, 38, 17, 18, 6, 7], (ii) tree-based learners [1], and (iii) meta-learners, which are model-agnostic and can be subdivided into direct learners (S-Learner and T-Learner) [25], and indirect learners (RA-Learner, DR-Learner, and X-Learner) [23, 11, 25]. For a comprehensive review, please refer to Curth et al. [10]. While these frameworks have demonstrated success, they predominantly address settings with a single treatment and a single outcome.
The study of composite treatment effects has been constrained by data scarcity arising from the exponential number of treatment combinations (for $K$ binary treatments, $2^{K}$ combinations) and the possibility that many combinations have insufficient or no data at all. In this context, Wang et al. [44] proposed an S-Learner-based approach, called deconfounder (DEC), that accommodates composite (binary and real) treatments for a single outcome. They leverage proxy variables and latent representations to capture interactions among multiple treatments and mitigate unmeasured confounding. Similarly, [49] developed the variational sample re-weighting (VSR) approach—an extension of S-Learner—to handle high-dimensional binary treatments (referred to as bundle treatments), using learned latent representations to de-correlate treatments from confounders. To address data scarcity, [33] proposed a data-augmentation strategy for composite treatment and single outcome settings, whereby $K$ ITE learners generate balanced datasets for each treatment before training a standard supervised deep learning model to predict potential outcomes. However, this method requires to train multiple models and is also limited to binary composite treatments.
@@ III BACKGROUND
Notations: Let $\mathbf{x}_{i}=(x_{i1},\dots,x_{ip})^{\top}$ denote the $p$-dimensional vector of pre-treatment features for patient $i$ ($i=1,\dots,N$). Let $\mathbf{t}_{i}=(t_{i1},\dots,t_{iK})^{\top}$ be the $K$-dimensional treatment vector for patient $i$, where $t_{ik}$ represents the $k$-th component of the potentially composite treatment. Let $\mathbf{y}_{i}=(y_{i1},\dots,y_{iM})^{\top}$ be the $M$-dimensional outcome vector for patient $i$, where $y_{im}$ is the $m$-th outcome. $y_{im}(\mathbf{t}_{i})$ denotes the potential outcome for patient $i$ on outcome $m$ if they were to receive treatment vector $\mathbf{t}_{i}$. The observed outcome is $y_{im}=y_{im}(\mathbf{t}_{i})$. The observational dataset is $\mathcal{D}=\{(\mathbf{x}_{i},\mathbf{t}_{i},\mathbf{y}_{i})\}_{i=1}^{N}$.
@@ IV METHODS
Motivation: Complex real-world scenarios often entail multiple treatments and multiple outcomes of interest. Existing research on ITE estimation, however, typically focuses on single treatment and single outcome problems. In the composite treatment and composite outcome setting, the exponential number of possible treatment combinations (e.g., $2^{K}$ for $K$ binary treatments) leads to data scarcity, where certain treatment combinations may have limited or no samples. Such data sparsity poses a significant challenge to reliably estimating ITE.
@@
Figure 2: Architecture of H-Learner: It comprises an embedding layer $e(t,y;\psi)$, a hypernetwork $h(e_{\mathbf{t},\mathbf{y}};\phi))$, and an ITE learner $f(\mathbf{X};\theta_{t,y})$. The hypernetwork $h$, conditioned on an embedding of treatment-outcome combination $e_{\mathbf{t},\mathbf{y}}$, generates the ITE learner’s $f$ weights $\theta_{t,y}$, enabling dynamic information sharing across treatments and outcomes.
@@
Figure 4: Comparative study of ITE learners for binary composite treatments and composite outcomes (shaded region represents one standard error over 10 repetitions). H-Learner does not need any assumptions about the interaction among the treatments and can handle composite outcomes, but SCP needs to know the causal structure among treatments and is limited to single outcomes.
Neural Architecture and Working: Figure 2 illustrates the key components of H-Learner, which include an embedding layer $e(t,y;\psi)$, a hypernetwork $h(e_{\mathbf{t},\mathbf{y}};\phi))$, and a target network (the ITE learner) $f(\mathbf{X};\theta_{t,y})$, where $\psi,\;\phi$ and $\theta$ are weights of embedding layer, hypernetwork and the ITE learner, respectively. The embedding layer learns representations for each unique combination of treatments and outcomes. It converts each outcome into a one-hot encoded vector and passes it through a linear layer to get the outcome embedding. To get treatment embedding, the treatment combination is passed through a linear layer, and the final embedding for the outcome and treatment combination is obtained by concatenating the respective embeddings. The embedding layer maps given causes and an effect to a meaningful representation, which is equivalent to learning embeddings for words in natural language processing [43].
@@
Training proceeds by feeding the treatments and outcomes into the embedding layer, whose outputs are passed to the hypernetwork. The hypernetwork produces weights for the target network, which processes the observed covariates to predict potential outcomes. We compute the prediction loss on factual observations (the data points for which the treatments and outcomes are known) and backpropagate the error. Since only the embedding layer and the hypernetwork parameters are trainable, the backpropagated gradients update those components to learn better embeddings for treatments and outcomes and improve the hypernetwork’s mapping from embeddings to target-network weights. Each training epoch thus leads to dynamic information sharing across the various treatments and outcomes, benefiting low-data configurations through mechanisms similar to transfer learning and multitasking [28].
@@ V-B Results
In this subsection, we compare H-Learner [^2] The finalised code will be available at https://github.com/jmdvinodjmd/HLearner under two scenarios: (1) arbitrary composite treatments and (2) binary composite treatments.